The area of the piston to the master cylinder in a hydraulic braking system of a car is 0.2 square inches. If a force of 5.9 lb is applied to this piston, what is the force (in pounds) on the brake pads of the wheels if the area of each pad is 1.3 square inches?
Answer #1

Here area of piston= 0.2 sq inch

And force applied to piston is 5.9lb

So firstly we find pressure .....so pressure = force / area

5.9/0.2=29.5 psi

Here we have each pad 1.3 square inch so basically there were four wheels in a car so each wheel have one pair of pad

Therefore for one wheel = 1.3 ×2 = 2.6 sq inch

So force = pressure × area

=29.5×2.6 =76.7 pound

As there are 4 wheels so ..= 4×76.7 = 306.8 pound